The length of a rectangle is 3 yd longer than its width.
If the perimeter of the rectangle is 34 yd, find its area

Answer by rapaljer(4671)

Let x = width
x+3= length

2W + 2L = P
2(x) + 2(x+3)=34
2x + 2x + 6=34
4x+6=34
4x=34-6
4x=28
x=7 yards = WIDTH
x+3=
7 + 3= 10 yards = LENGTH

Check:
Perimeter = 2*7+2*10
=14+20
=34 yards. It checks!!

Now, the area of the rectangle = W*L= 3*10= 30 square yards!

For additional help with WORD PROBLEMS, please take a look at my own website. To see it click on my tutor name "rapaljer" anywhere in algebra.com. Look for the link to my website there. From my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time", click on "Basic Algebra," and then "Chapter 1," which is my chapter introducing the basics of algebra, including Equation Solving and Word Problems. Look in "Sections 1.09 and 1.10".

See also the solutions of many of these exercises in the "MATH IN LIVING COLOR" pages, where the problems are solved in COLOR. My own students and others have been unanimous in telling me that my own explanations are easier to understand than the expensive textbooks that are written for mathematicians and publishers!! MY EXPLANATIONS WERE WRITTEN FOR STUDENTS WHO HAVE A HARD TIME UNDERSTANDING MATH!! And they are all FREE!!

I hope you like my website!!! If you do, please recommend it to your family and friends!! At least check out the JOKE PAGE--all safe jokes for the little ears in your family!!

Merry Christmas!!!!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus